OpenStudy (anonymous):

solve 5-2(3z+2)=-4(1-2z)

OpenStudy (anonymous):



OpenStudy (compassionate):

Hi, 5-2(3z+2)=-4(1-2z) [Distribute -2 and -4] 5 - 6z + 4 = -4 + 8z [Add like terms] 9 - 6z = -4 + 8z [Transpose] -6z - 8z = -9 - 4 [add like terms] -14z = -13 [Divide] z = 0.9

OpenStudy (anonymous):

I agree